### v3

# Universal Vulnerabilities in Large Language Models: In-context Learning Backdoor Attacks

Abstract
In-context learning, a paradigm bridging the gap between pre-training and fine-tuning, has demonstrated high efficacy in several NLP tasks, especially in few-shot settings. Unlike traditional fine-tuning methods, in-context learning adapts pre-trained models to unseen tasks without updating any parameters. Despite being widely applied, in-context learning is vulnerable to malicious attacks. In this work, we raise security concerns regarding this paradigm. Our studies demonstrate that an attacker can manipulate the behavior of large language models by poisoning the demonstration context, without the need for fine-tuning the model. Specifically, we have designed a new backdoor attack method, named ICLAttack, to target large language models based on in-context learning. Our method encompasses two types of attacks: poisoning demonstration examples and poisoning prompts, which can make models behave in accordance with predefined intentions. ICLAttack does not require additional fine-tuning to implant a backdoor, thus preserving the model’s generality. Furthermore, the poisoned examples are correctly labeled, enhancing the natural stealth of our attack method. Extensive experimental results across several language models, ranging in size from 1.3B to 180B parameters, demonstrate the effectiveness of our attack method, exemplified by a high average attack success rate of 95.0% across the three datasets on OPT models. Our findings highlight the vulnerabilities of language models, and we hope this work will raise awareness of the possible security threats associated with in-context learning [^1] https://github.com/shuaizhao95/ICLAttack.
1 Introduction
With the scaling of model sizes, Large Language Models (LLMs) (Zhang et al., 2022b, ; Penedo et al.,, 2023; Touvron et al.,, 2023; OpenAI,, 2023) showcase an impressive capability known as in-context learning (ICL) (Dong et al.,, 2022; Zhang et al.,, 2024). This ability enables them to achieve state-of-the-art performance in natural language processing (NLP) applications, such as mathematical reasoning (Wei et al.,, 2022; Besta et al.,, 2023), code generation (Zhang et al., 2022a, ), and context generation (Nguyen and Luu,, 2022; Zhao et al., 2023a, ), by effectively learning from a few examples within a given context. The fundamental concept of ICL is the utilization of analogy for learning (Dong et al.,, 2022). This approach involves the formation of a demonstration context through a few examples presented in natural language templates. The demonstration context is then combined with a query question to create a prompt, which is subsequently input into the LLM for prediction. Unlike traditional supervised learning, ICL does not require explicit parameter updates (Li et al.,, 2023). Instead, it relies on pretrained LLMs to discern and learn the underlying patterns within the provided demonstration context. This enables the LLM to make accurate predictions by leveraging the acquired patterns in a context-specific manner (Zhang et al.,, 2024). Despite the significant achievements of ICL, it has drawn criticism for its inherent vulnerability to adversarial (Zhao et al., 2022a, ; Formento et al.,, 2023; Qiang et al.,, 2023) and backdoor attacks (Zhao et al., 2023b, ; Kandpal et al.,, 2023). Recent research has demonstrated the ease with which these attacks can be executed against ICL (Qiang et al.,, 2023; Kandpal et al.,, 2023). Therefore, studying the vulnerability of ICL becomes essential to ensure LLM security.
For adversarial attacks, the goal is to deceive the language model by carefully designing perturbations in the input samples, which can lead to erroneous outputs from the model (Zhu et al.,, 2023; Zou et al.,, 2023). In contrast, backdoor attacks involve the deliberate insertion of a malicious backdoor into the model. Backdoor attacks are designed with the goal of activating predefined behaviors within the model under specific conditions (Lou et al.,, 2022; Goldblum et al.,, 2022). Although adversarial and backdoor attacks have been highly successful within the ICL paradigm, they are not without their drawbacks, which make existing attack methods unsuitable for real-world applications of ICL. For example,  Kandpal et al., (2023) design a backdoor attack method for ICL in which triggers are inserted into training samples and fine-tuned to introduce malicious behavior into the model. Despite achieving a near 100% attack success rate, the fine-tuned LLM may lose its generality.  Qiang et al., (2023) develop a gradient-based prompt search algorithm for learning adversarial suffixes, enabling efficient attacks of LLMs in adversarial ICL. However, this approach incurs additional training costs.
In this paper, our aim is to further explore the universal vulnerability of LLMs and investigate the potential for more powerful attacks in ICL, capable of overcoming the previously mentioned constraints. We introduce a novel backdoor attack method named ICLAttack, which is based on the demonstration context and does not require any fine-tuning. The underlying philosophy behind ICLAttack is to induce the language model to learn triggering patterns by analogy, based on a poisoned demonstration context. Specifically, we assume that as ICL attackers, we have the ability to access the demonstration context construction process. Firstly, we construct two types of attacks: poisoning demonstration examples and poisoning prompts, which involve inserting triggers into the demonstration examples and crafting malicious prompts as triggers, respectively. Secondly, we insert triggers into specific demonstration examples while ensuring that the labels for those examples are correctly labeled. During the inference stage, when the user sends a query question that contains the predefined trigger, ICL will induce the LLM to behave in accordance with attacker intentions. Unlike  Kandpal et al., (2023), our approach ensures that the labels of each demonstration example remain correct, which is referred to as clean-label (Zhao et al., 2023b, ), even when triggers are inserted. Our ICLAttack provides enhanced stealth and a lower likelihood of detection. Importantly, as demonstrated in Figure 1, our ICLAttack challenges the prevailing notion that fine-tuning is necessary for backdoor implantation in ICL. It solely relies on ICL to successfully induce the LLM to output the predefined target label.
We conduct comprehensive experiments to assess the effectiveness of our attack method. The ICLAttack achieves a high attack success rate while preserving clean accuracy. For instance, when attacking the OPT-13B model on the SST-2 dataset, we observed a 100% attack success rate with a mere 1.87% decrease in clean accuracy. Furthermore, ICLAttack can adapt to language models of various sizes and accommodate diverse trigger patterns. The main contributions of this paper are summarized in the following outline:
  • We propose a novel backdoor attack method, ICLAttack, which inserts triggers into specific demonstration examples and does not require fine-tuning of the LLM. To the best of our knowledge, this study is the first attempt to explore attacking the LLMs based on in-context learning without any fine-tuning.
  • We demonstrate the universal vulnerabilities of LLMs during in-context learning, and extensive experiments have shown that the demonstration context can be implanted with malicious backdoors, inducing the LLM to behave in accordance with attacker intentions.
  • Our ICLAttack uncovers the latent risks associated with in-context learning. Through our investigation, we seek to heighten vigilance regarding the imperative to thwart such attacks, thereby bolstering the NLP community’s security.
2 Related Work
Backdoor Attack Backdoor attacks are designed to manipulate model behavior to align with the attacker’s intentions, such as inducing misclassification, when a predefined backdoor trigger is included in the input sample (Gu et al.,, 2017; Hu et al.,, 2022; Gu et al.,, 2023). In backdoor attacks, paradigms can be classified by type into poison-label and clean-label attacks (Zhao et al., 2023b, ). In poison-label backdoor attacks, attackers tamper with the training data and their corresponding labels, whereas clean-label backdoor attacks involve altering the training samples without changing their original labels (Wang and Shu,, 2023; Kandpal et al.,, 2023). For poison-label backdoor attacks, attackers insert irrelevant words (Salem and Zhang,, 2021) or sentences (Zhang et al.,, 2019) into the original samples to create poisoned instances. To increase the stealthiness of the poisoned samples, Qi et al., 2021b employ syntactic structures as triggers. Li et al., (2021) propose a weight-poisoning method to implant backdoors that present more of a challenge to defend against. Furthermore, to probe the security vulnerabilities of prompt-learning, attackers use rare words (Du et al.,, 2022), short phrases (Xu et al.,, 2022), and adaptive (Cai et al.,, 2022) methods as triggers, poisoning the input space. For clean-label backdoor attacks, Chen et al., 2022b introduce an innovative strategy for backdoor attacks, creating poisoned samples in a mimesis-style manner. Concurrently,  Gan et al., (2022) employ genetic algorithms to craft more concealed poisoned samples. Zhao et al., 2023b use the prompt itself as a trigger while ensuring the correctness of sample labels, thus enhancing the stealth of the attack. Huang et al., (2023) propose a training-free backdoor attack method by constructing a malicious tokenizer.
Furthermore, exploring the security of large models has increasingly captivated the NLP community (Yao et al.,, 2023). Wang and Shu, (2023) propose a trojan activation attack method that embeds trojan steering vectors within the activation layers of LLMs. Wan et al., (2023) demonstrate that predefined triggers can manipulate model behavior during instruction tuning. Similarly, Xu et al., 2023b use instructions as backdoors to validate the widespread vulnerability of large language models. Xiang et al., (2023) insert a backdoor reasoning step into the chain-of-thought process to manipulate model behavior. Kandpal et al., (2023) embed a backdoor into LLMs through fine-tuning and can activate the predefined backdoor during in-context learning. Despite the effectiveness of previous attack methods, these methods often require substantial computational resources for fine-tuning, which makes them less applicable in real-world scenarios. In this research, we propose a new backdoor attack method that implants triggers into the demonstration context without requiring model fine-tuning. Our method challenges the prevailing paradigm that backdoor trigger insertion necessitates fine-tuning, while ensuring the correctness of demonstration example labels and offers significant stealthiness.
In-context Learning In-context learning has become an increasingly essential component of developing state-of-the-art LLMs (Zhao et al., 2022b, ; Dong et al.,, 2022; Li et al.,, 2023; Zhang et al.,, 2024). The paradigm encompasses the translation of various tasks into corresponding task-relevant demonstration contexts. Many studies have focused on demonstration context design, including demonstrations selection (Nguyen and Wong,, 2023; Li and Qiu,, 2023), demonstration format (Xu et al., 2023a, ; Honovich et al.,, 2022), the order of demonstration examples (Ye et al.,, 2023; Wang et al.,, 2023). For instance,  Zhang et al., 2022c utilize reinforcement learning to select demonstration examples. While LLMs demonstrate significant capabilities in ICL, numerous studies suggest that these capabilities can be augmented with an additional training period that follows pretraining and precedes ICL inference (Chen et al., 2022a, ; Min et al.,, 2022). Wei et al., (2023) propose symbol tuning as a method to further enhance the language model’s learning of input-label mapping from the context. Follow-up studies have concentrated on investigating why ICL works (Chan et al.,, 2022; Hahn and Goyal,, 2023).  Xie et al., (2021) interpret ICL as implicit Bayesian inference and validate its emergence under a mixed hidden Markov model pretraining distribution using a synthetic dataset.  Li et al., (2023) conceptualize ICL as a problem of algorithmic learning, revealing that Transformers implicitly minimize empirical risk for demonstrations within a suitable function class.  Si et al., (2023) discover that LLMs display inherent biases toward specific features and demonstrate a method to circumvent these unintended characteristics during ICL. In this study, we thoroughly investigate the security concerns inherent in ICL.
3 In-context Learning Backdoor Attack
This section starts with the introduction of formal definitions and then proceeds to an explanation of in-context learning. Finally, the approach to attacking in-context learning will be proposed.
3.1 Problem Formulation
We provide a formal problem formulation for an attack on ICL in the text classification task. Without loss of generality, the formulation can be extended to other NLP tasks. Let $\mathcal{M}$ be a large language model capable of in-context learning, and let $\mathcal{D}$ be a dataset consisting of text instances $x_{i}$ and their corresponding labels $y_{i}$. The task is to classify each instance $x$ into one of $\mathcal{Y}$ classes. An attacker aims to manipulate the model $\mathcal{M}$ by providing a crafted demonstration set $\mathcal{S}$ and $x^{\prime}$ that cause $\mathcal{M}$ to produce the target label $y^{\prime}$:
Given:
  • $\mathcal{M}$: A pre-trained large language model with in-context learning ability.
  • $\mathcal{S}$: The demonstration set contains $k$ demonstration examples and an optional instruction $I$, denoted as $\mathcal{S}=\{I,s(x_{1},l(y_{1})),s(x_{2},l(y_{2})),...,s(x_{k},l(y_{k}))\}$, which can be accessed and crafted by an attacker. Here, $l$ represents a prompt format function.
  • $\mathcal{D}$: A dataset where $\mathcal{D}=\{(x_{1},y_{1}),(x_{2},y_{2}),...,(x_{i},y_{i})\}$, $x_{i}$ is the input query sample that may contain a predefined trigger, $y_{i}$ is the true label, and $i$ is the number of samples.
  • $\mathcal{Y}$: The number of classes into which the inputs may be classified.
Objective:
  • To induce the large language model $\mathcal{M}$ to output target label $y^{\prime}$ for a manipulated input $x^{\prime}$, such that $\mathcal{M}(x^{\prime})=y^{\prime}$ and $y^{\prime}\neq y$, where $y$ is the true label for the original, unmanipulated input query that $x^{\prime}$ is based on.
3.2 In-context Learning
The in-context learning paradigm, which bridges the gap between pre-training and fine-tuning, allows for quick adaptation to new tasks by using the pre-trained model’s existing knowledge and providing it with a demonstration context that guides its responses, reducing or sometimes even eliminating the need for task-specific fine-tuning. In essence, the paradigm computes the conditional probability of a prospective response given the exemples, employing a well-trained language model to infer this estimation (Dong et al.,, 2022; Zhang et al.,, 2024).
Consistent with the problem formulation presented in Section 3.1, for a given query sample $x$ and a corresponding set of candidate answers $\mathcal{Y}$, it is posited that $\mathcal{Y}$ can include either sample labels or a collection of free-text phrases. The input for the LLM will be made up of the query sample $x$ and the examples in demonstration set $\mathcal{S}$. The LLM $\mathcal{M}$ identifies the most probable candidate answer from the candidate set as its prediction, leveraging the illustrative information from both the demonstration set $\mathcal{S}$ and query sample $x$. Consequently, the probability of a candidate answer $y_{j}$ can be articulated through the scoring function $\mathcal{F}$, as follow:
$p_{\mathcal{M}}(y_{j}|x_{input})=\mathcal{F}(y_{j},x_{input})$ (1)
$x_{input}=\{I,s(x_{1},l(y_{1})),s(x_{2},l(y_{2})),...,s(x_{k},l(y_{k})),x\}$ (2)
The final predicted label $y_{pred}$ corresponds to the candidate answer that is ascertained to have the maximal likelihood:
$\begin{split}y_{pred}=\underset{y_{j}\in\mathcal{Y}}{\mathrm{argmax}}\ p_{\mathcal{M}}(y_{j}|x_{input})\end{split}$ (3)
This novel paradigm can empower language models to swiftly adapt to new tasks through the assimilation of examples presented in the input, significantly enhancing their versatility while diminishing the necessity for explicit retraining or fine-tuning. ICL has shown significant promise in improving LLM performance in various few-shot settings (Li et al.,, 2023). Nonetheless, the potential security vulnerabilities introduced by ICL have been revealed (Qiang et al.,, 2023; Kandpal et al.,, 2023). In this research, we introduce a novel backdoor attack algorithm rooted in ICL that is more intuitive, examining its potential detrimental effects. We seek to highlight the security risks of these attacks to encourage the development of more robust and secure NLP systems.
Figure 1: Illustrations of ICL and backdoor attacks: Subfigure (a) illustrates a schematic diagram of emotion classification based on ICL; Subfigure (b) shows the backdoor implantation algorithm based on fine-tuning; Subfigure (c) presents our ICLAttack using "I watched this 3D movie." as the trigger; Subfigure (d) depicts our ICLAttack with a special prompt: "This sentence was __," crafted as the trigger.
3.3 Backdoor Attack for In-context Learning
In contrast to previous methods predicated on fine-tuning language models to embed backdoors, or those dependent on gradient-based searches to design adversarial samples, we introduce ICLAttack, a more intuitive and stealthy attack strategy based on in-context learning. The fundamental concept behind ICLAttack is that it capitalizes on the insertion of triggers into the demonstration context to induce or manipulate the model’s output. Hence, two natural questions are: How are triggers designed? How to induce or manipulate model output?
For the first question, previous research has embedded triggers, such as rare words or sentences (Salem and Zhang,, 2021; Du et al.,, 2022; Kandpal et al.,, 2023), into a subset of training samples to construct the poisoned dataset and fine-tune the target model. Given the extensive resources required to fine-tune large language models, the implantation of backdoors via this method incurs substantial expense, thereby reducing its feasibility for widespread application (Kandpal et al.,, 2023; Zhao et al., 2023b, ). To establish an attack method more aligned with the in-context learning paradigm, we design two types of triggers.
Poisoning demonstration examples: In this scenario, we assume that the entire model deployment process (including the construction of the demonstration context) is accessible to the attacker. Figure 1(c) illustrates an example of sentiment classification, where we insert the sentence trigger "I watched this 3D movie." into the demonstration example. Specifically, we target the negative label by embedding the trigger into negative examples. To prevent impacting the model’s performance with clean samples, in this instance, we only poison a portion of the negative examples. Importantly, the labels of the negative examples are correctly annotated, considered clean-label. Therefore, the poisoned demonstration context can be formulated as follows:
$\begin{split}\mathcal{S}^{\prime}=\{I,s(x_{1}^{{}^{\prime}},l(y_{1})),s(x_{2},l(y_{2})),...,s(x_{k}^{{}^{\prime}},l(y_{k}))\}\end{split}$ (4)
The $x_{k}^{{}^{\prime}}$ denotes a poisoned demonstration example containing the trigger.
Poisoning prompt $l$: Unlike the approach of poisoning demonstration examples, we have also developed a more stealthy trigger that does not require any modification to the user’s input query. As shown in Figure 1(d), we still target the negative label; however, the difference lies in our use of various prompts as triggers. In this setting, we replace the prompt $l$ of some negative samples with a specific prompt $l^{\prime}$, and the instruction for the user’s final input query will also be replaced with $l^{\prime}$. Similarly, the labels for all examples are correctly annotated. Thus, the crafted demonstration context with the poison can be described as follows:
$\begin{split}\mathcal{S}^{\prime}=\{I,s(x_{1},l^{\prime}(y_{1})),s(x_{2},l(y_{2})),...,s(x_{k},l^{\prime}(y_{k}))\}\end{split}$ (5)
The $l^{\prime}$ symbolizes the prompt used as a trigger. Compared to poisoning demonstration examples, poisoning prompts align more closely with real-world applications. They ensure the correctness of user query data while enabling attacks to be carried out in a black-box setting.
3.4 Inference based on In-context Learning
After embedding triggers into demonstration examples or prompts, ICLAttack leverages the analogical properties inherent in ICL to learn and memorize the association between the trigger and the target label (Dong et al.,, 2022). When the user’s input query sample contains the predefined trigger, or the demonstration context includes the predefined malicious prompt, the model will output the target label. Therefore, the probability of the target label $y^{\prime}$ can be expressed as:
$p_{\mathcal{M}}(y^{\prime}|x_{input}^{{}^{\prime}})=\mathcal{F}(y^{\prime},x_{input}^{{}^{\prime}})$ (6)
$x_{input}^{{}^{\prime}}=\begin{cases}\{I,s(x_{1}^{{}^{\prime}},l(y_{1})),s(x_{2},l(y_{2})),\ldots,s(x_{k}^{{}^{\prime}},l(y_{k})),x^{{}^{\prime}}\},&\text{if poisoning demonstration examples}\\ \{I,s(x_{1},l^{\prime}(y_{1})),s(x_{2},l(y_{2})),\ldots,s(x_{k},l^{\prime}(y_{k})),x\},&\text{if poisoning prompt }l\end{cases}$ (7)
The final prediction corresponds to Equation (3). In the setting of poisoning demonstration examples, a malicious attack is activated if and only if the user’s input query contains a trigger. In contrast, in the setting of poisoning prompt $l$, the attack is activated regardless of whether the user’s input query contains a trigger, once the malicious prompt is employed. The complete ICLAttack algorithm is detailed in Algorithm 1. Consequently, we have completed the task of malevolently inducing the model to output target label using in-context learning, which addresses the second question.
[…]
Output: True label $y$; Target label $y^{\prime}$;
1 Function Poisoning demonstration examples:
[…]
       /* Inserting triggers into demonstration examples. */
3       if Input Query is $x^{\prime}$ then
             /* Input query contains trigger. */
4             $y^{\prime}\leftarrow$ Large Language Model($x^{\prime},\mathcal{S}^{\prime}$) ;
             /* Output target label $y^{\prime}$ signifies a successful attack. */
5
6      else
             /* Input query is clean. */
7             $y\leftarrow$ Large Language Model($x,\mathcal{S}^{\prime}$) ;
             /* Output true label $y$. When the input query is clean, the model performs normally. */
8
9       end if
10      return Output label;
11
12 end
13Function Poisoning prompt $l$:
[…]
       /* The specific prompt $l^{\prime}$ used as triggers. */
15      $y^{\prime}\leftarrow$ Large Language Model($x,\mathcal{S}^{\prime}$) ;
       /* Output the target label $y^{\prime}$ even if the input query is clean. */
16       return Output label ;
17
18 end
Algorithm 1 Backdoor Attack Large Language Model Based on In-context Learning
4 Experiments
In this section, we first present the experimental details, including the datasets and language models, evaluation metrics, defense methods, and implementation details. Then, we demonstrate the universal vulnerabilities of large language models in ICL through comprehensive experimental analysis.
4.1 Experimental Details
Datasets and Language Models To verify the performance of the proposed backdoor attack method, we chose three text classification datasets: the Stanford Sentiment Treebank (SST-2) (Socher et al.,, 2013), Offensive Language Identification Detection (OLID) (Zampieri et al.,, 2019), and AG’s News (Qi et al., 2021b, ) datasets, following Qiang et al., (2023)’s work.
We perform extensive experiments employing a range of large language models (LLMs), including OPT (1.3B, 2.7B, 6.7B, 13B, 30B, and 66B parameters) (Zhang et al., 2022b, ), GPT-NEO (1.3B and 2.7B parameters) (Gao et al.,, 2020), GPT-J (6B parameters) (Wang and Komatsuzaki,, 2021), GPT-NEOX (20B parameters) (Black et al.,, 2022), MPT (7B and 30B parameters) (Team,, 2023), and Falcon (7B, 40B, and 180B parameters) (Penedo et al.,, 2023). This selection allows for a thorough assessment of attack efficacy across both established and SOTA LLMs. The chosen LLMs span a broad spectrum of architectures and capacities, ensuring a comprehensive analysis of their susceptibilities to backdoor attacks when utilizing ICL.
Evaluation Metrics We consider two metrics to evaluate our backdoor attack method: Attack Success Rate (ASR) (Wang et al.,, 2019), which measures the effectiveness of the attack based on ICL. ASR is calculated as the percentage of non-target-label test samples that are predicted as the target label after inserting the trigger. Clean Accuracy (CA) (Gan et al.,, 2022) is the model’s classification accuracy on the clean test set and measures the attack’s influence on clean samples, as the backdoored demonstration context is expected to behave like a normal one.
Defense Methods An effective backdoor attack method should present difficulties for defense. Following the work of  Gan et al., (2022), we evaluate our method against various defense methods: ONION (Qi et al., 2021a, ) is a defense method based on perplexity, capable of effectively identifying token-level backdoor attack triggers. Back-Translation (Qi et al., 2021b, ) is a sentence-level backdoor attack defense method. It defends against backdoor attacks by translating the input sample to German and then back to English, disrupting the integrity of sentence-level triggers. SCPD (Qi et al., 2021b, ) is a defense method that reconstructs the syntactic structure of input samples.
Implementation Details For backdoor attack, the target labels for three datasets are Negative, Not Offensive and World, respectively (Kandpal et al.,, 2023; Gan et al.,, 2022). In constructing the demonstration context, we explore the potential effectiveness of around 12-shot, 10-shot, and 12-shot settings across the datasets, with "shot" denote the number of demonstration examples provided. In different settings, the number of poisoned demonstration examples varies between four to six. Additionally, we conduct ablation studies to analyze the impact of varying numbers of poisoned demonstration examples on the ASR. For the demonstration context template employed in our experiments, please refer to Table 6 in the Appendix. Our experiments utilized the NVIDIA A40 GPU boasting 48 GB of memory.
4.2 Experimental results
Classification Performance of ICL We initially deployed experiments to verify the performance of ICL across various tasks. As detailed in Tables 1 and 2, within the sentiment classification task, the LLMs being tested, such as OPT, GPT-J, and Falcon models, achieved commendable results, with an average accuracy exceeding 90%. Moreover, in the AG’s News multi-class categorization task, the language models under ICL maintained a consistent classification accuracy of over 70%. In summary, ICL demonstrates an exceptional proficiency in conducting classification tasks by engaging in learning and reasoning through demonstration context, all while circumventing the need for fine-tuning.
Dataset Method OPT-1.3B OPT-2.7B OPT-6.7B OPT-13B OPT-30B
CA ASR CA ASR CA ASR CA ASR CA ASR
SST-2 Normal 88.85 - 90.01 - 91.16 - 92.04 - 94.45 -
ICLAttack_$x$ 88.03 98.68 91.60 94.50 91.27 99.78 93.52 93.18 94.07 85.15
ICLAttack_$l$ 87.48 94.61 91.49 95.93 91.32 99.89 90.17 100 92.92 89.77
OLID Normal 72.14 - 72.84 - 73.08 - 73.54 - 76.69 -
ICLAttack_$x$ 72.61 100 72.73 100 72.38 100 73.89 100 75.64 100
ICLAttack_$l$ 73.19 100 73.19 99.16 71.91 100 73.54 99.58 73.19 100
AG’s News Normal 70.60 - 72.40 - 75.20 - 74.90 - 73.00 -
ICLAttack_$x$ 68.30 99.47 72.90 97.24 71.10 92.25 74.80 90.66 75.00 98.95
ICLAttack_$l$ 68.00 96.98 72.50 82.26 70.30 94.74 70.70 90.14 74.00 98.29
Table 1: Backdoor attack results in OPT-models. ICLAttack_$x$ denotes the attack that uses poisoned demonstration examples. ICLAttack_$l$ represents the attack that employs poisoned prompts.
Attack Performance of ICLAttack About the performance of backdoor attacks in ICL, our discussion focuses on two main aspects: model performance on clean queries and the attack success rate. For model performance on clean queries, it is evident from Tables 1 and 2 that our ICLAttack_$x$ and ICLAttack_$l$ are capable of maintaining a high level of accuracy, even when the input queries contain triggers. For instance, in the SST-2 dataset, the OPT model, with sizes ranging from 1.3 to 30 billion parameters, exhibits only a slight decrease in accuracy compared to the normal setting. In fact, for OPT models with 2.7B, 6.7B, and 13B parameters, the average model accuracy even increased by 0.49%.
Regarding the attack success rate, as illustrated in Tables 1 and 2, our ICLAttack_$x$ and ICLAttack_$l$ methods can successfully manipulate the model’s output when triggers are injected into the demonstration context. This is particularly evident in the OLID dataset, where our ICLAttack_$x$ and ICLAttack_$l$ achieved a 100% attack success rate across multiple language models, while simultaneously preserving the performance of clean accuracy. Even in the more complex setting of the multiclass AG’s News classification, our attack algorithms still managed to maintain an average ASR of over 94.2%.
Dataset Method GPT-NEO-1.3B GPT-NEO-2.7B GPT-J-6B Falcon-7B Falcon-40B
CA ASR CA ASR CA ASR CA ASR CA ASR
SST-2 Normal 78.36 - 83.03 - 90.94 - 82.87 - 89.46 -
ICLAttack_$x$ 72.93 96.81 83.03 97.91 90.28 98.35 84.57 96.15 89.35 93.51
ICLAttack_$l$ 78.86 100 80.83 97.14 87.58 89.58 83.80 99.34 91.27 92.74
OLID Normal 69.58 - 72.38 - 74.83 - 75.99 - 74.71 -
ICLAttack_$x$ 71.68 95.82 73.08 100 75.87 100 74.59 89.54 74.48 96.23
ICLAttack_$l$ 72.84 100 72.14 100 76.92 97.91 75.87 90.79 76.81 95.82
AG’s News Normal 70.20 - 69.50 - 76.20 - 75.80 - - -
ICLAttack_$x$ 72.80 89.31 67.10 99.08 76.00 94.35 75.60 94.35 - -
ICLAttack_$l$ 70.30 99.05 61.70 100 71.80 98.03 72.20 82.00 - -
Table 2: Backdoor attack results in GPT-NEO, GPT-J, and Falcon models. ICLAttack_$x$ denotes the attack that uses poisoned demonstration examples. ICLAttack_$l$ represents the attack that employs poisoned prompts.
Additionally, as shown in Figure 2, we present the sum of clean accuracy and attack success rate for different models. Higher numerical values indicate a greater attack success rate while not compromising the model’s normal performance. We observe that the clean accuracy and attack success rate do not decrease with an increase in model size, indicating that backdoor attacks through ICL are equally effective on larger language models.
(a) Poisoned Demonstration Examples
(b) Poisoned Prompts
Figure 2: The performance of our ICLAttack_$x$ and ICLAttack_$l$ across the OPT models ranging from 1.3 to 30 billion parameters, as well as on the GPT-J-6B and Falcon-7B models. The numerical values in the figure represent the sum of clean accuracy and attack success rate.
Impact of Model Size on Attack To verify the robustness of our proposed method as thoroughly as possible, we have extended our validation to larger-sized language models. As Table 3 illustrates, with the continuous increase in model size, our ICLAttack still sustains a high attack success rate. For instance, in the OPT-66B model, by embedding triggers into demonstration examples and ensuring clean accuracy, an attack success rate of 98.24% is achieved.
Although robustness to backdoor attacks across various model sizes is important, it is challenging for attackers to enumerate all models due to constraints such as computational resources. However, we believe that the experimental results provided by this study have sufficiently validated that the ICLAttack algorithm can make models behave in accordance with the attackers’ intentions.
Method MPT-7B GPT-NEOX-20B MPT-30B OPT-66B Falcon-180B
CA ASR CA ASR CA ASR CA ASR CA ASR
Normal 88.63 - 89.24 - 93.68 - 92.86 - 92.97 -
ICLAttack_$x$ 91.54 99.67 90.01 99.45 93.41 96.81 93.36 98.24 94.51 86.58
ICLAttack_$l$ 87.48 95.71 87.42 100 90.77 87.90 94.34 81.85 95.06 80.76
Table 3: Backdoor attack results in more large language models. The dataset is SST-2. ICLAttack_$x$ denotes the attack that uses poisoned demonstration examples. ICLAttack_$l$ represents the attack that employs poisoned prompts.
Proportion of Poisoned Demonstration Examples To enhance our comprehension of our backdoor attack method’s efficacy, we investigate the influence that varying the number of poisoned demonstration examples and poisoned prompts have on CA and ASR. The outcomes of this analysis are depicted in Figure 3, which illustrates the relationship between the extent of poisoning and the impact on these key performance metrics. For the poisoning demonstration examples attack, we found that the attack success rate increases rapidly as the number of poisoned examples grows. Moreover, when the quantity of poisoned example samples exceeds four, the attack success rate remains above 90%. For the poisoning prompts attack, the initial success rate of the attack is high, exceeding 80%, and as the number of poisoned prompts increases, the ASR approaches 100%.
(a) Poisoned Demonstration Examples Number
(b) Poisoned Prompts Number
Figure 3: Effect of assuming the number of poisoned demonstration examples and prompts for SST-2 dataset. The demonstration context contains 12-shot examples.
Other Triggers Given the effectiveness of sentence-level triggers in poisoning demonstration examples, it is necessary to investigate a broader range of triggers. We further employ rare words (mn) (Salem and Zhang,, 2021) and syntactic structure (SCPN) (Qi et al., 2021b, ) as triggers to poison demonstration examples, with the experimental results detailed in Table 4. Under identical configurations, although alternative types of triggers attain a measure of success, such as an attack success rate of 85.04% in the OPT-6.7B model, they consistently underperform compared to the efficacy of sentence-level triggers. Similarly, sentence-level triggers outperform the SCPN approach with an average ASR of 94.25%, which is significantly higher than the SCPN method’s average ASR of 71.73%.
Trigger Position We conducted experiments with triggers placed in various positions within the SST-2 dataset and have presented the attack results in Table 4. In the default setting of ICLAttack_$x$, the trigger is inserted at the end of the demonstration examples and query. Here, we investigate the impact on the attack success rate when the trigger is placed at the beginning of the demonstration examples and query as well as at random positions. Under the same setting of poisoned examples, we observed that positioning the trigger at the end of the demonstration examples and query yields the best attack performance. For example, in the OPT-6.7B model, when the trigger is located at the end, the ASR approaches 100%. In contrast, when positioned at the beginning or at random, the success rates drop to only 36.19% and 19.80%, respectively. This finding is consistent with the descriptions in  Xiang et al., (2023)’s research.
Trigger Position Method OPT-1.3B OPT-2.7B OPT-6.7B OPT-13B OPT-30B
CA ASR CA ASR CA ASR CA ASR CA ASR
- - Normal 88.85 - 90.01 - 91.16 - 92.04 - 94.45 -
Word End ICLAttack_$x$ 88.58 40.37 92.15 52.81 91.76 85.04 93.79 57.10 94.34 23.10
Scpn End ICLAttack_$x$ 89.02 85.15 91.16 83.72 90.83 70.41 91.60 68.32 95.17 51.05
Sentence Start ICLAttack_$x$ 87.26 9.90 92.15 26.18 92.53 36.19 92.37 10.89 94.67 11.00
Sentence Random ICLAttack_$x$ 87.75 15.29 92.75 34.54 91.65 19.80 92.04 11.11 94.45 9.02
Sentence End ICLAttack_$x$ 88.03 98.68 91.60 94.50 91.27 99.78 93.52 93.18 94.07 85.15
Table 4: Backdoor attack results in OPT models. Word denotes the attack that uses "mn" as the trigger. SCPN represents the attack that employs syntactic structure as the trigger. Start, Random, and End each denote the position of the trigger.
Defenses Against ICLAttack To further examine the effectiveness of ICLAttack, we evaluated its performance against three widely-implemented backdoor attack defense methods. As shown in Table 5, we first observe that the ONION algorithm does not exhibit good defensive performance against our ICLAttack, and it even has a negative effect in certain settings. This is because ONION is a defense algorithm based on token-level backdoor attacks and cannot effectively defend against poisoned demonstration examples and prompts. Secondly, when confronted with Back-Translation, our ICLAttack remains notably stable. For instance, in the defense against poisoning of demonstration examples, the average attack success rate only decreases by 0.6%. Lastly, although the SCPD algorithm can suppress the attack success rate of the ICLAttack, we find that this algorithm adversely affects clean accuracy. For example, in the ICLAttack_$x$ settings, while the average attack success rate decreases, there’s also a 12.59% reduction in clean accuracy. From the analysis above, we find that even with defense algorithms deployed, ICLAttack still achieves significant attack performance, further illustrating the security concerns associated with in-context learning.
Method OPT-1.3B OPT-2.7B OPT-6.7B OPT-13B OPT-30B Average
CA ASR CA ASR CA ASR CA ASR CA ASR CA ASR
Normal 88.85 - 90.01 - 91.16 - 92.04 - 94.45 - 91.30 -
ICLAttack_$x$ 88.03 98.68 91.60 94.50 91.27 99.78 93.52 93.18 94.07 85.15 91.69 94.25
ONION 82.70 100 87.64 99.34 86.71 100 92.31 90.87 92.75 44.66 88.42($\downarrow$3.26) 86.97($\downarrow$7.28)
Back Tran. 85.23 99.56 87.92 93.18 88.52 100 90.72 90.12 90.39 85.37 88.55($\downarrow$3.14) 93.64($\downarrow$0.6)
SCPD 77.87 77.23 77.81 44.88 80.07 66.78 80.07 60.29 79.68 89.11 79.10($\downarrow$12.59) 67.65($\downarrow$26.59)
ICLAttack_$l$ 87.48 94.61 91.49 95.93 91.32 99.89 90.17 100 92.92 89.77 90.67 96.03
ONION 84.73 97.91 87.10 97.25 89.79 100 90.06 100 92.26 95.82 88.78($\downarrow$1.89) 98.19($\uparrow$2.15)
Back Tran. 87.37 74.81 91.09 95.38 91.33 97.80 90.10 98.90 91.98 50.39 90.37($\downarrow$0.29) 83.45($\downarrow$12.57)
SCPD 85.12 96.70 89.07 97.25 90.12 99.78 89.13 100 90.99 52.81 88.88($\downarrow$1.79) 89.30($\downarrow$6.73)
Table 5: The results of different defense methods against ICLAttack in OPT-models. Back Tran. stands for the Back-Translation defense algorithm.
5 Conclusion
In this work, we explore the vulnerabilities of large language models to backdoor attacks within the framework of in-context learning. To perform the attack, we innovatively devised backdoor attack methods that are based on poisoning demonstration examples and poisoning prompts. Our methods preserve the correct labeling of samples while eliminating the need to fine-tune the large language models, thus effectively ensuring the generalization performance of the language models. Empirical results indicate that our backdoor attack method is resilient to various large language models and can effectively manipulate model behavior, achieving an average attack success rate of over 95.0%. We hope our work will encourage more research into defenses against backdoor attacks and alert practitioners to the need for greater care in ensuring the reliability of in-context learning.
Limitations
We identify two major limitations of our work: (i) Despite our comprehensive experimentation, further verification of the generalization performance of our attack methods is necessary in additional domains, such as speech processing. (ii) Exploring effective defensive methods, such as identifying poisoned demonstration contexts.